The director of a summer day camp estimates that 100 children will join if the camp fee is $250, but for each $20 decrease in the fee, ten more children will enroll. A.)Determine the linear equation that will represent the number of children who will enroll at a given fee. B.)Approximately how many students will enroll if the camp fee is $180? Round to the nearest child. C.) How many students will enrol if the camp is free? Round to the nearest child.
Answer by stanbon(48568)

The director of a summer day camp estimates that 100 children will join if the camp fee is $250, but for each $20 decrease in the fee, ten more children will enroll.
A.)Determine the linear equation that will represent the number of children who will enroll at a given fee.
N(x) = 100+10(250-x)/20
-----------------------------------------------------

B.)Approximately how many students will enroll if the camp fee is $180? Round to the nearest child.
N(180) = 100+10(250-180)/20
N(180) = 100 + 10(7/2)
N(180) = 135
------------------------------

C.) How many students will enroll if the camp is free? Round to the nearest child.
N(0) = 100 + 10(250-0)/20
N(0) = 100 + 10(12.5)
N(0) = 100 + 125
N(0) = 225
=====================
Cheers,
Stan H.